shay opened up an savings account with 600 dollars. At a yearly interest rate of 5% that is compounded monthly, how much interest will she earn in six months
Answer by stanbon(75887)

shay opened up a savings account with 600 dollars.
At a yearly interest rate of 5% that is compounded monthly,
how much interest will she earn in six months?
---------------------------------
A(t) = p(1+(r/n))^(nt)
-----
A(1/2) = 600(1+(0.05/0.5))^(0.5*0.05)
-----
A(1/2) = 600(1+0.1)^(0.025)
A(1/2) = 600*1.1^0.025
A(1/2) = 601.43
----------------------
Interest was $1.43
===========================
Cheers,
stan H.